OpenStudy (anonymous):

find the solution for the following linear system: y=1/2x-3 y=3/2x-1

OpenStudy (anonymous):

is there a particular way u wanna solve it with?

OpenStudy (anonymous):

y=1/2x-3 y=3/2x-1 1/2x-3=3/2x-1 -1/2x -1/2x subtract 1/2x from both sides -3 = x -1 +1 +1 add 1 to both sides x = -2 plug in into either equation y=1/2x-3 y=1/2(-2)-3 y = -4 solution (-2,-4)